OpenStudy (anonymous):

Hi! Can anybody help me with this solution problem? 9x+3y=15 3x=5-y

OpenStudy (anonymous):

By substitution or addition?

OpenStudy (anonymous):

Sorry Thats addition

OpenStudy (anonymous):

Okay. First lets rewrite the second equation as 3x + y = 5. If we multiply the second equation by -3, we get -9x - 3y = -15. Adding to two together 9x + 3y = 15 -9x - 3y = -15 We see that everything equals zero. This means both equations are linearly dependent. In other words, they are the same equation.